# OWT: A Foundational Organ-Wise Tokenization Framework for Medical Imaging

### v1

@@ Abstract
Recent advances in representation learning often rely on holistic, black-box embeddings that entangle multiple semantic components, limiting interpretability and generalization. These issues are especially critical in medical imaging. To address these limitations, we propose an Organ-Wise Tokenization (OWT) framework with a Token Group-based Reconstruction (TGR) training paradigm. Unlike conventional approaches that produce holistic features, OWT explicitly disentangles an image into separable token groups, each corresponding to a distinct organ or semantic entity. Our design ensures each token group encapsulates organ-specific information, boosting interpretability, generalization, and efficiency while allowing fine-grained control in downstream tasks. Experiments on CT and MRI datasets demonstrate the effectiveness of OWT in not only achieving strong image reconstruction and segmentation performance, but also enabling novel semantic-level generation and retrieval applications that are out of reach for standard holistic embedding methods. These findings underscore the potential of OWT as a foundational framework for semantically disentangled representation learning, offering broad scalability and applicability to real-world medical imaging scenarios and beyond.  [^1] Code is available at https://github.com/SifanSong/OWT
@@ 1 Introduction
Recent years have witnessed significant advances in representation learning of visual data for both natural and medical images. The typical representation learning, driven by self-supervised and task-guided training frameworks, is classified as holistic embedding-based representation learning since it learns embeddings that capture all information for an entire image [54, 22, 58] (Fig.1-a). Similar trends appear in the medical imaging domain, where holistic representations have also been extensively explored for tasks like disease diagnosis[49], organ segmentation [46, 5, 25], and anatomical structure analysis [41, 59]. Although the holistic embeddings leveraged by these studies have been proven effective, they remain globally entangled and mixed in a black-box latent space, combining information about multiple organs and background. This fusion decreases efficiency and poses significant challenges for organ-focused analyses, disease diagnosis, or radiological interpretation, where fine-grained interpretability and a semantically targeted focus are critically important [48, 6, 58].
@@
In parallel, disentangled representation learning (DRL) has been studied to decompose data into semantically meaningful factors of variation, with each factor ideally corresponding to an independent aspect of the underlying data [58] (Fig.1-b). By isolating latent factors (e.g., shape, color, orientation) in natural images [4, 12, 61, 64] or modality attributes in medical images [44, 8], DRL methods aim to improve explainability and enable controllable feature manipulation. However, most conventional DRL approaches rely on purely data-driven or perturbation-driven strategies, requiring extensive modeling assumptions or specialized regularizers to ensure each factor corresponds to a semantically coherent concept [38, 58]. Consequently, existing DRL approaches face notable challenges, including limited alignment with real-world semantics, weak spatial correspondence, and poor scalability efficiency.
@@
Despite these successes, existing holistic embedding–based strategies and DRL approaches still have notable gaps, particularly in medical imaging, where high-level interpretability and organ-specific analysis are crucial (e.g., anatomical analysis and surgical planning). Specifically, existing methods suffer from (a) a lack of interpretability and generalizability, as multiple semantic cues are often merged without alignment to real-world concepts, making it difficult to isolate relevant structures from confounding factors (e.g., background variations); and (b) limited efficiency and controllability, since these approaches process entire or masked representations rather than focusing only on organ-level features, restricting the ability to selectively manipulate or analyze distinct semantic factors.
These limitations motivate us to revisit the design of representation learning, aiming for explicit organ- (or part-) level tokenization to tackle them. In this paper, we propose a foundational Organ-Wise Tokenization (OWT) framework and a novel Token Group-based Reconstruction (TGR) training paradigm. Specifically, in addition to a standard encoder-decoder architecture, such as Masked Autoencoders (MAE) [27], our OWT framework incorporates two key novel modules, Organ Collector and Adaptive Holistic Embedding Restorer (AHER), inspired by prior structures for extracting semantic features [47, 18]. The Organ Collector integrates holistic embeddings into semantically meaningful token groups, while AHER reconstructs holistic representations from any number of token groups. Together, these modules enable a structured transition from holistic $\rightarrow$ semantic $\rightarrow$ holistic embeddings, facilitating both organ-specific processing and reconstruction. Meanwhile, our proposed TGR training paradigm ensures that the extracted semantic embeddings are fully functional as Semantically Disentangled Token Groups (SDTGs) by explicitly correlating each token group to its corresponding organ. In this case, each SDTG encapsulates only organ-specific information and is capable of being utilized independently or in any combination with other SDTGs as practically needed (Fig. 1-c).
Therefore, our design offers key advantages by addressing the limitations of existing approaches: (i) improved interpretability and generalizability, as each token group is explicitly associated with a specific organ or semantic content, enabling the isolation of target structures while minimizing confounding factors; and (ii) enhanced efficiency and controllability, as OWT processes only the relevant token groups for a given task, reducing computational load and allowing for semantic-based, fine-grained manipulation and adaptation without introducing unrelated features.
@@
We evaluate the performance of the organ-wise embeddings (i.e., SDTGs) extracted by OWT across multiple downstream tasks (such as image reconstruction, segmentation and retrieval). Across multiple modalities (CT, two MRI sequences, and a facial dataset), our method achieves better and more robust results compared to baselines (i.e., MAE [27] and dedicated semantic segmentation networks [63, 42, 9, 5, 51]) and is able to accomplish tasks that these traditional methods cannot, such as semantic-level generation, retrieval and embedding combination. This proves the concept of our foundational semantically disentangled tokenization framework, enabling us to bring more potential solutions for real-world applications.
@@
  • We propose a foundational framework and a novel training paradigm, OWT and TGR, which encourage the network to encapsulate organ-level features into semantically disentangled token groups (SDTGs) rather than entangled embeddings with a black-box nature.
  • The proposed SDTGs are highly condensed with rich information on each target organ, which increases the interpretability and generalization of latent embeddings.
  • The token groups can be used separately or in combination and are processed with improved efficiency and controllability for downstream tasks and further analysis.
  • Experiments demonstrate that our method offers significant scalability. It can adaptively integrate semantic information based on various requirements and incorporate other modalities (e.g., text encoding), highlighting its potential for broader applications beyond medical imaging.
2 Related Work
2.1 Holistic Embedding-based Representation Learning
Holistic embedding-based approaches map entire 2D/3D inputs into a single latent representation, aiming to capture global image features [54, 22, 58]. These methods typically follow either self-supervised or task-guided strategies. In self-supervised learning, contrastive techniques like SimCLR [10], SimSiam [11], and BYOL [24] enforce consistency between augmented views, while predictive approaches such as CPC [43] and RotNet [23] reconstruct missing or permuted portions. Recently, masked image modeling (e.g., MAE [27] and BEiT [3]) has emerged as an effective approach by requiring models to infer masked regions. Meanwhile, task-guided frameworks focus on extracting features that directly benefit a target task, typically resulting in holistic embeddings. Many classic neural networks, such as ResNet [26], DenseNet [29], and Vision Transformer [16], and their variants incorporate classification, regression, or segmentation objectives.
In medical imaging, holistic representations have been widely used for self-supervised feature learning in CT [13, 32], MRI [7, 33], and X-ray [52, 60]. Additionally, classical supervised pipelines (e.g., U-Net [46] and its variants [63, 9, 5, 25]) rely on a global bottleneck to embed entire images or volumes, demonstrating strong performance in organ segmentation and disease detection. Although such approaches achieve notable success, they often encode multiple semantic factors, such as multiple organs and the surrounding background, into an entangled representation with limited interpretability and fine-grained control [48, 6, 58].
2.2 Disentangled Representation Learning
In contrast to holistic embedding-based approaches, various disentangled representation learning (DRL) studies have pursued a dimension-wise strategy for two main objectives. First, to derive more robust holistic embeddings with improved generalizability [14, 1, 15, 56, 20, 50, 58], and second, to explicitly decompose images into latent factors, each corresponding to distinct and semantically meaningful attributes, in order to improve explainability [4, 12, 61, 64].
In medical imaging, the motivations for DRL are especially strong: improved interpretability, controllability, and robustness to confounding factors are crucial in clinical settings [37, 55]. When transitioning DRL to medical imaging, researchers seek to decouple anatomical properties (e.g., organ size or shape) [57, 30], modality characteristics (e.g., CT and MRI sequences) [44, 8], and pathological factors (e.g., tumor presence or severity) [36, 28], aiming to isolate distinct clinical features.
Despite notable progress, data- and perturbation-driven DRL approaches can be sensitive to training objectives, architecture choices, and assumptions about independent factors [38, 56, 58]. In existing studies, disentangled dimensions may not align precisely with real anatomical or pathological concepts, and spatial correspondence often remains weak when images are globally encoded. Consequently, there has been a growing emphasis on methods that combine disentanglement principles with controlled grouping or tokenization schemes to encapsulate semantically meaningful subsets of features.
@@ 3 Method
In this paper, we present a novel OWT framework and a Token Group-based Reconstruction (TGR) training paradigm, which embeds each organ into its corresponding semantically disentangled token group (SDTG). In this case, these organ-specific embeddings can be utilized independently or in combination for downstream tasks.
3.1 Overall OWT Framework
As illustrated in Fig. 2, our proposed OWT framework comprises five components: an encoder, Organ Collector, Token Group Encoder, Adaptive Holistic Embedding Restorer (AHER), and a decoder. The encoder maps input images to holistic embeddings, while the Organ Collector and TGR enable the target-organ features to be extracted and stored in specific embeddings (i.e., SDTGs). In TGR, we randomly select a subset of the predefined token groups. The retained token groups are fed into the Token Group Encoder to integrate both inter-organ and intra-organ information. Next, AHER enables any number of retained token groups to be converted back into a holistic embedding. Finally, the decoder reconstructs an image of the same size as the input.
The encoder and decoder in OWT follow standard holistic representation learning architectures, so they can be Transformer-based or (fused) CNN-based, as in 2D/3D MAE [27, 19] or 2D/3D VQGAN [17, 19]. This design choice allows the framework to be applied to either 2D or 3D images with minimal modification. For a 3D input volume processed by a Transformer-based encoder, let $I\in\mathbb{R}^{D\times H\times W\times C}$. The encoder ($\mathbf{Enc}$) produces a holistic embedding $X_{H}=\mathbf{Enc}(I)\in\mathbb{R}^{(dhw)\times c_{e}}$, where $D$, $H$, $W$ and $C$ represent the depth, height, width, and channels of the volume, respectively, while $d$, $h$ and $w$ are the corresponding embedded dimensions. The $(dhw)$ represents the number of tokens processed by the Transformer blocks, with each token having an embedding dimension of $c_{e}$. This holistic representation is then fed into subsequent modules for more fine-grained organ-wise processing.
@@ 3.2 Organ Collector
To effectively extract semantically meaningful features from the holistic embedding $X_{H}$, we introduce a novel Organ Collector module. A key component of this process is the token-wise attention matrix $A_{t}$, which dynamically assigns importance to different spatial locations or tokens, ensuring that each token group captures the most relevant semantic features.
To achieve this, let $X_{H}\in\mathbb{R}^{(dhw)\times c_{e}}$ be reshaped to $\mathbb{R}^{c_{e}\times d\times h\times w}$ before being fed into the Organ Collector (Fig. 2). We then employ a token-wise attention mechanism, defined as:
@@
where $N_{g}$ denotes the number of semantic-based tokens, and both $\alpha$ and $\gamma$ are linear layers. Concretely, $\alpha$ projects the $(c_{e}\times d)$-dimensional input to $N_{g}$, then applies Flatten and Softmax to derive a token-wise attention matrix. This attention matrix is multiplied by $\gamma(X_{H})$ (also flattened) to obtain $X_{G}\in\mathbb{R}^{N_{g}\times c_{e}}$.
Abdomen1k Holistic-based Rec. 1 Token Group Rec.
L2 $\downarrow$ LPIPS $\downarrow$ SSIM-3D $\uparrow$ L2 $\downarrow$ LPIPS $\downarrow$ SSIM-3D $\uparrow$
MAE [27] 6.74e-04 0.0684 0.9688 - - -
MAE/${}_{\text{LPIPS}}$ 6.19e-04 0.0403 0.9722 - - -
OWT 4.16e-04 0.0390 0.9734 3.31e-04 0.0073 0.9795
OWT/${}_{\text{4slices}}$ 3.69e-04 0.0371 0.9755 2.68e-04 0.0069 0.9838
AutoPet Holistic-based Rec. 1 Token Group Rec.
L2 $\downarrow$ LPIPS $\downarrow$ SSIM-3D $\uparrow$ L2 $\downarrow$ LPIPS $\downarrow$ SSIM-3D $\uparrow$
MAE [27] 2.86e-04 0.0452 0.9799 - - -
MAE/${}_{\text{LPIPS}}$ 2.77e-04 0.0255 0.9843 - - -
OWT 1.49e-04 0.0232 0.9881 3.32e-04 0.0071 0.9778
OWT/${}_{\text{4slices}}$ 1.13e-04 0.0179 0.9908 2.68e-04 0.0063 0.9827
Table 1: Performance of image reconstruction on CT datasets.
Abdomen1k Holistic-based Rec. 1 Token Group Rec.
L2 $\downarrow$ LPIPS $\downarrow$ SSIM-3D $\uparrow$ L2 $\downarrow$ LPIPS $\downarrow$ SSIM-3D $\uparrow$
MAE [27] 1.37e-03 0.0465 0.9686 - - -
MAE/${}_{\text{LPIPS}}$ 1.36e-03 0.0389 0.9690 - - -
OWT 7.65e-04 0.0256 0.9851 5.96e-04 0.0088 0.9762
OWT/${}_{\text{4slices}}$ 5.81e-04 0.0214 0.9885 5.04e-04 0.0081 0.9801
PreArtery Holistic-based Rec. 1 Token Group Rec.
L2 $\downarrow$ LPIPS $\downarrow$ SSIM-3D $\uparrow$ L2 $\downarrow$ LPIPS $\downarrow$ SSIM-3D $\uparrow$
MAE [27] 1.67e-03 0.0438 0.9701 - - -
MAE/${}_{\text{LPIPS}}$ 1.74e-03 0.0375 0.9672 - - -
OWT 8.66e-04 0.0248 0.9851 5.29e-04 0.0089 0.9759
OWT/${}_{\text{4slices}}$ 7.07e-04 0.0218 0.9876 4.47e-04 0.0082 0.9801
Table 2: Performance of image reconstruction on MRI datasets.
@@ 3.3 Token Group-based Reconstruction (TGR)
The TGR training paradigm provides direct supervision to token groups, ensuring their semantic disentanglement. By enforcing accurate organ-wise feature extraction and reconstruction through the token group-level loss $\mathcal{L}_{T}$ (Fig. 3-b), TGR preserves both spatial and textural consistency. Consequently, the extracted semantic tokens serve as semantically disentangled token groups (SDTGs). Through TGR, the Organ Collector adaptively extracts organ-specific information into its corresponding SDTG, assigning each group to a particular organ.
@@
Definition of Token Group. Given a dataset $D$, we have its ground-truth annotations for $g$ organs. The remaining regions are treated as background and assigned to a single additional token group. Thus, we define a total of $(g+1)$ token groups. Let each group contain $f$ tokens, making the total number of tokens $N_{g}=(g+1)\times f$. For example, in Fig. 2, $X_{0}$, $X_{1}$, and $X_{2}$ represent the token groups that store liver, kidney, and background information, respectively.
@@
Random Selection Procedure. To manage these token groups so that they are responsible for each organ during training, we adopt a two-stage random selection procedure. First, randomly determine the number of retained groups: Let $\Omega\sim\mathcal{U}(0,1)$ be a uniform random variable on $[0,1]$. With $(g+1)$ total Token Groups, we define $\tilde{g}=\bigl{\lfloor}(g+1)\cdot\Omega\bigr{\rfloor}$, resulting in an integer $\tilde{g}\in\{0,1,\dots,g\}$. Second, randomly select $\tilde{g}$ groups: From the full set of $(g+1)$ token groups, we then randomly pick exactly $\tilde{g}$ groups to retain and discard the rest. We set each group to contain $f$ tokens so that each separable and independent token group is in $\mathbb{R}^{{f}\times c_{e}}$. We concatenate the $\tilde{g}$ selected groups along the token dimension, producing the final retained tokens $\tilde{X}_{G}\in\mathbb{R}^{{\tilde{N}_{g}}\times c_{e}}$ (Fig 2), where the final retained token count is $\tilde{N}_{g}=\tilde{g}\times f$.
@@
During TGR, for each input image, we remain only areas according to randomly retain token groups based on organ annotations (Fig. 3-b), treating the processed image as the target $I_{target}$ for the token group-based reconstruction loss $\mathcal{L}_{T}$. This explicit supervision encourages a precise organ-level focus, preserving each organ’s unique spatial and textural details in the retained token groups (i.e., SDTGs). Meanwhile, any randomly excluded token groups are discarded at this stage, ensuring that only the retained groups contribute to reconstruction and thereby enhancing semantic flexibility and computational efficiency. As illustrated in Fig. 3, TGR departs from conventional holistic embedding-based reconstruction methods that rely on the latent space for the entire image [2, 27, 19, 17] (Fig. 3-a). By selectively preserving certain token groups, our approach avoids fusing diverse organ features and does not require masked tokens, resulting in improved interpretability, efficiency, and controllability during both reconstruction and subsequent tasks.
@@
After that, we employ a Token Group Encoder ($\mathbf{Enc}_{G}$) composed of six multi-head linear self-attention blocks [31]. Let $\tilde{X}^{\prime}_{G}=\mathbf{Enc}_{G}(\tilde{X}_{G})\in\mathbb{R}^{\tilde{N}_{g}\times c_{e}}$ represent the further contextual modeling that captures inter- and intra-organ relationships of the retained token groups.

### v2

@@ Abstract
Recent advances in representation learning often rely on holistic embeddings that entangle multiple semantic components, limiting interpretability and generalization. These issues are especially critical in medical imaging, where downstream tasks depend on anatomically interpretable features. To address these limitations, we propose an Organ-Wise Tokenization (OWT) framework with a Token Group-based Reconstruction (TGR) training paradigm. Unlike conventional approaches, OWT explicitly disentangles an image into separable token groups, each corresponding to a distinct organ or semantic entity. Our design ensures each token group encapsulates organ-specific information, boosting interpretability, generalization, and efficiency while enabling fine-grained control for targeted clinical applications. Experiments on CT and MRI datasets demonstrate OWT’s power: it not only achieves strong performance on standard tasks like image reconstruction and segmentation, but also unlocks novel, high-impact clinical capabilities including organ-specific tumor identification, organ-level retrieval and semantic-level generation, without requiring any additional training. These findings underscore the potential of OWT as a foundational framework for semantically disentangled representation learning, offering broad scalability and a new perspective on how representations can be leveraged.
@@ 1 Introduction
Representation learning has advanced rapidly in recent years, with growing impact across both natural and medical imaging domains, enabling more effective feature extraction and improved performance across a wide range of downstream tasks. Building on this progress, a variety of self-supervised and task-guided training frameworks have been widely adopted, typically relying on holistic embeddings that capture aggregated information for an entire image without explicitly separating semantically distinct components. [voulodimos2018deep, geirhos2020shortcut, wang2024disentangled] (Fig.1-a). Similar trends appear in the medical imaging domain, where holistic representations have also been extensively explored for tasks like disease diagnosis [shen2017deep], organ segmentation [ronneberger2015u, cao2022swin, hatamizadeh2021swin], and anatomical structure analysis [mortada2023segmentation, weston2019automated]. Although holistic embeddings have shown strong performance, they remain globally mixed within a latent space, entangling information across multiple organs and background regions without explicit semantic separation. This mixed representation poses significant challenges for where fine-grained interpretability and a semantically targeted focus are critically important [schutte2021using, cetin2023attri, wang2024disentangled].
@@
On the other hand, disentangled representation learning (DRL) aims to decompose data into semantically meaningful factors, each capturing an independent aspect of the underlying variation [wang2024disentangled] (Fig.1-b). DRL isolates latent factors (e.g., shape, color, orientation) in natural images [burgess2018understanding, chen2016infogan, xiao2017dna, zhu2018visual] or modality attributes in medical images [ouyang2021representation, chartsias2019disentangled], thereby improving explainability and enabling controllable feature manipulation. However, most existing DRL approaches rely on purely data-driven or perturbation-driven strategies, requiring extensive modeling assumptions or specialized regularizers to ensure each factor corresponds to a semantically coherent concept [locatello2019challenging, wang2024disentangled]. Consequently, existing DRL approaches suffer from limited alignment with real-world semantics, weak spatial correspondence, and poor scalability (More details on Related Work are provided in Appendix A).
@@
While DRL offers a promising direction, both holistic embeddings and existing disentanglement approaches remain limited in medical imaging scenarios, where organ-level interpretability and analysis are important (e.g., anatomical analysis and surgical planning). These limitations manifest in two key aspects: (a) a lack of interpretability and generalizability, as multiple semantic cues are often entangled without alignment to real-world concepts, making it difficult to isolate relevant structures from confounding factors (e.g., background); and (b) limited efficiency and controllability, as these approaches operate on fused representations rather than isolated organ-level features, restricting the ability to selectively manipulate or analyze distinct semantic factors. These limitations motivate the development of our novel representation learning framework that explicitly tokenizes images at the organ or semantic level.
In this paper, we propose a foundational Organ-Wise Tokenization (OWT) framework and a novel Token Group-based Reconstruction (TGR) training paradigm (Fig. 1-c). The OWT framework extends a standard encoder-decoder architecture, such as Masked Autoencoders (MAE) or VQGAN, by incorporating two specialized modules: Organ Collector and the Adaptive Holistic Embedding Restorer (AHER), which are inspired by prior work on semantic tokenization [ryoo2021tokenlearner, esser2021taming, fayyaz2022adaptive].
The Organ Collector integrates holistic embeddings into semantically meaningful Token Group (TG), while AHER enables to reconstruct holistic representations from any number of token groups. This design enables a transition pathway from holistic $\rightarrow$ semantic $\rightarrow$ holistic embeddings, facilitating both organ-specific processing and reconstruction. To ensure that the extracted semantic embeddings function as true Semantically Disentangled Token Groups (SDTGs), the proposed TGR training paradigm explicitly enforces correspondence between each token group and its target organ. These components constitute a unified framework that encodes organ-specific semantics into modular and reusable TGs, enabling task-adaptive control and restricting computation to relevant semantic regions without additional training, thereby enhancing interpretability and eliminating redundant processing.
@@
We evaluate the organ-wise embeddings on downstream tasks (e.g., image reconstruction and segmentation). Across multiple modalities (CT, two MRI sequences, and a facial dataset), our method consistently outperforms strong baselines, including VAE [kingma2013auto], MAE [he2022masked], and state-of-the-art task-specific segmentation architectures such as UNet++ [zhou2018unet++], AttU_Net [oktay2018attention], TransUNet [chen2021transunet], Swin-UNet [cao2022swin], and CMUNet [tang2023cmu]. Beyond these general tasks, OWT unlocks capabilities that conventional models cannot provide: organ-level tumor identification and retrieval, semantic-level generation, and flexible combination of organ embeddings to adapt to heterogeneous downstream requirements. This demonstrates that the semantic disentanglement enabled by our OWT framework can be leveraged for broader real-world applications, offering a new perspective on how representations can be constructed and utilized.
@@
  • We introduce OWT, the first organ-wise foundational framework that, paired with a novel TGR training paradigm, encapsulates organ-specific semantics into semantically disentangled token groups (SDTGs) within a single training process.
  • The learned SDTGs are highly condensed yet rich in information for each target organ, enhancing the interpretability and generalizability of the latent embeddings.
  • The SDTGs can be selectively combined and reused, yielding improved efficiency and fine-grained controllability. This unlocks unique organ-level downstream tasks, such as organ-specific retrieval and identification, which are highly demanded in the medical imaging domain.
  • Our experiments demonstrate significant scalability, as OWT adaptively disentangles semantic information across diverse modalities and domains. This includes a real-world facial dataset, highlighting its potential for applications beyond medical imaging.
@@ 2 Method
We propose Organ-Wise Tokenization (OWT) with Token Group-based Reconstruction (TGR): OWT encodes each organ into a semantically disentangled token group (SDTG) that can be used alone or composed for downstream tasks.
Overall, the proposed OWT framework comprises five components (Fig. 2): an encoder, Organ Collector, Token Group Encoder, Adaptive Holistic Embedding Restorer (AHER), and a decoder. The encoder maps input images to holistic embeddings, while the Organ Collector and TGR enable the target-organ features to be extracted and stored in specific embeddings (i.e., SDTGs). In TGR, we randomly select a subset of the predefined token groups. The retained token groups are fed into the Token Group Encoder to integrate both inter-organ and intra-organ information. Next, AHER enables any number of retained token groups to be converted back into a holistic embedding. Finally, the decoder reconstructs an image of the same size as the input.
2.1 Encoder and Decoder in OWT
The encoder and decoder in OWT follow standard holistic representation learning architectures, so they can be Transformer-based or (fused) CNN-based, as in 2D/3D MAE [he2022masked, feichtenhofer2022masked] or 2D/3D VQGAN [esser2021taming, feichtenhofer2022masked]. This design choice allows the framework to be applied to either 2D or 3D images with minimal modification. For a 3D input volume processed by a Transformer-based encoder, let $I\in\mathbb{R}^{D\times H\times W\times C}$. The encoder ($\mathbf{Enc}$) produces a holistic embedding $X_{H}=\mathbf{Enc}(I)\in\mathbb{R}^{(dhw)\times c_{e}}$, where $D$, $H$, $W$ and $C$ represent the depth, height, width, and channels of the volume, respectively, while $d$, $h$ and $w$ are the corresponding embedded dimensions. The $(dhw)$ represents the number of tokens processed by the Transformer blocks, with each token having an embedding dimension of $c_{e}$. This holistic representation is then fed into subsequent modules for more fine-grained organ-wise processing.
@@ 2.2 Organ Collector
To effectively extract semantically meaningful features from holistic embedding $X_{H}$, the Organ Collector module dynamically assigns importance to different spatial locations or tokens using a token-wise attention matrix $A_{t}$.
Let $X_{H}\in\mathbb{R}^{(dhw)\times c_{e}}$ be reshaped to $\mathbb{R}^{c_{e}\times d\times h\times w}$ before being fed into the Organ Collector (Fig. 2). We then employ a token-wise attention mechanism, defined as:
@@
where $N_{g}$ denotes the number of semantic-based tokens, and both $\alpha$ and $\gamma$ are linear layers. Concretely, $\alpha$ projects the $(c_{e}\times d)$-dimensional input to $N_{g}$, then applies Flatten and Softmax to derive a token-wise attention matrix. This attention matrix is multiplied by $\gamma(X_{H})$ (also flattened and transposed) to obtain $X_{G}\in\mathbb{R}^{N_{g}\times c_{e}}$. This ensures that each token group captures the most relevant semantic features.
@@ 2.3 Token Group-based Reconstruction (TGR)
The TGR training paradigm provides direct supervision of semantic disentanglement to token groups. By enforcing accurate organ-wise feature extraction and reconstruction through the token group-level loss $\mathcal{L}_{T}$ (Fig. 3-b), TGR preserves both spatial and textural consistency. Consequently, the extracted semantic tokens serve as semantically disentangled token groups (SDTGs). Through TGR, the Organ Collector adaptively extracts organ-specific information into its corresponding SDTG, assigning each token group to a particular organ.
@@
Definition of Token Group. Given a dataset $D$, we have its ground-truth annotations for $g$ semantic group (such as organs in medical image). The remaining regions are treated as background and assigned to a single additional token group. Thus, we define a total of $(g+1)$ token groups. Let each group contain $f$ tokens, making the total number of tokens $N_{g}=(g+1)\times f$. For example, in Fig. 2, $X_{0}$, $X_{1}$, and $X_{2}$ represent the token groups that store liver, kidney, and background information, respectively.
@@
Random Selection Procedure. To establish and preserve the correspondence between token groups and their target organs during training, we apply a two-stage random selection. First, randomly determine the number of retained groups: Let $\Omega\sim\mathcal{U}(0,1)$ be a uniform random variable on $[0,1]$. With $(g+1)$ total Token Groups, we define $\tilde{g}=\bigl\lfloor(g+1)\cdot\Omega\bigr\rfloor$, resulting in an integer $\tilde{g}\in\{0,1,\dots,g\}$. Second, randomly select $\tilde{g}$ groups: From the full set of $(g+1)$ token groups, we then randomly pick exactly $\tilde{g}$ groups to retain and discard the rest. We set each group to contain $f$ tokens so that each separable and independent token group is in $\mathbb{R}^{{f}\times c_{e}}$. We concatenate the $\tilde{g}$ selected groups along the token dimension, producing the final retained tokens $\tilde{X}_{G}\in\mathbb{R}^{{\tilde{N}_{g}}\times c_{e}}$ (Fig 2), where the final retained token count is $\tilde{N}_{g}=\tilde{g}\times f$.
@@
During TGR, for each input image, we remain only areas according to randomly retain token groups based on organ annotations (Fig. 3-b), treating the processed image as the target $I_{target}$ for the token group-based reconstruction loss $\mathcal{L}_{T}$. This explicit supervision encourages a precise organ-level focus, preserving each organ’s unique spatial and textural details in the retained token groups (i.e., SDTGs). Meanwhile, any randomly excluded token groups are discarded at this stage, ensuring that only the retained groups contribute to reconstruction and thereby enhancing semantic flexibility and computational efficiency. As illustrated in Fig. 3, TGR departs from conventional holistic embedding-based reconstruction methods that rely on the latent space for the entire image [bank2023autoencoders, he2022masked, feichtenhofer2022masked, esser2021taming] (Fig. 3-a). By selectively preserving certain token groups, our approach avoids fusing diverse organ features and does not require masked tokens, resulting in improved interpretability, efficiency, and controllability during both reconstruction and subsequent tasks.
@@
After that, we employ a Token Group Encoder ($\mathbf{Enc}_{G}$) composed of six multi-head linear self-attention blocks [katharopoulos2020transformers]. Let $\tilde{X}^{\prime}_{G}=\mathbf{Enc}_{G}(\tilde{X}_{G})\in\mathbb{R}^{\tilde{N}_{g}\times c_{e}}$ represent the further contextual modeling that captures inter- and intra-organ relationships of the retained token groups.